Verify that the differential equation (y^2+yz)dx+(xz+z^2)dy+(y^2-xy)dz=0 is integrable and find its primitive


1
Expert's answer
2021-06-14T15:33:17-0400

The necessary and sufficient condition for iintegrability is


"X\cdot curl X=0"



"X=(y^2+yz,xz+z^2,y^2-xy)" so that



"\nabla\times X=\begin{vmatrix}
\vec i & \vec j & \vec k \\
 {\partial\over \partial x} & {\partial\over \partial y} & {\partial\over \partial z}
\\
 y^2+yz & xz+z^2 & y^2-xy
\end{vmatrix}"




"=(2y-x-x-2z)\vec i+(y+y)\vec j + (z-2y-z)\vec k"




"=(2y-2x-2z)\vec i+(2y)\vec j + (-2y)\vec k"




"X\cdot (\nabla\times X)=2y^3-2xy^2-2y^2z+2y^2z"




"-2xyz-2yz^2+2xyz+2yz^2-2y^3+2xy^2=0"

Thus the given equation is integrable.


Solve by Inspection


"y(y+z)dx+z(x+z)dy+y(y-x)dz=0"

Or


"y(y+z)dx+y(y+z)dz-y(y+z)dz"




"+z(x+z)dy+y(x+z)dy-y(x+z)dy"




"+y(y-x)dz=0"

Or


"y(y+z)d(x+z)+(y+z)(x+z)dy"




"-ydz(y+z-y+x)-y(x+z)dy=0"

Or



"y(y+z)d(x+z)+(y+z)(x+z)dy-y(x+z)d(y+z)=0"




"{d(x+z) \over x+z}+{dy\over y}-{d(y+z) \over y+z}=0"

The complete primitive is given as


"y(x+z)=c(y+z)"